OpenStudy (rockycatsrule):

Which ordered pair in the form (a, b) is a solution of this equation? 7a – 5b = 28 A. (0, 4) B. (–2, –3) C. (–3, –2) D. (4, 0)

OpenStudy (rushwr):

tell me your answer !

OpenStudy (rockycatsrule):

maybe ether a or d

OpenStudy (rushwr):

The answer is D

OpenStudy (rockycatsrule):

k

OpenStudy (rushwr):

A is not possible because if a = 0 then the whole part (7a) becomes zero According to the equation 7a - 5b = 28 if 7a becomes zero 28 = -5b then b should have a negative value. But in the answer A u have a positive value for b with the zero for a so that's wrong. D is correct because when u substitute it to the equation 7*4 - (5*0) = 28 28 - 0 =28 Then u get the answer right!

OpenStudy (rockycatsrule):

For which points is the sum of their x- and y-coordinates greater than 4 [for example, (5, 2) since 5 + 2 = 7, or (–3, 9) since –3 + 9 = 6]? A. E, X B. Z, J C. X, L D. D, L

OpenStudy (rushwr):

What are you answers ?

OpenStudy (rockycatsrule):

c,or b, maybe a??????????/

OpenStudy (rushwr):

LOL okai, the answer is a

OpenStudy (rockycatsrule):

k lol

OpenStudy (rockycatsrule):

Which point is the graph of the ordered pair (–4, 2)? A. point C B. point A C. point D D. point B answers r b, d, or c

TheSmartOne (thesmartone):

it would be better if you made a new post for your question :)

OpenStudy (rockycatsrule):

ok

OpenStudy (rushwr):

C - the x co-ordinate = -4 the y co-ordinate = +5 That sums up to +1 (5-4 = 1) D - the x co- ordinate = -9 The y co-ordinate = -9 that sums up to 0

OpenStudy (rushwr):

E - the x co-ordinate = +5 the y co-ordinate = 0 that sums up to +5 which is greater than 4 so that's in

OpenStudy (rushwr):

j- the x co-ordinate = -9 the y co=ordinate = 0 Sums up to -9

OpenStudy (rushwr):

L - the x- co-ordinate = +9 the y co=ordinate = -9 sums upto 0

OpenStudy (rushwr):

p - the x co-ordinate = 0 the y co-ordinate = -9 sums up to -9

OpenStudy (rushwr):

x - the x co-ordinate = 0 the y co-ordinate = +5 that sums up to +5 which is greater than 4 so an answer

OpenStudy (rushwr):

Z - the x co-ordinate = 10 the y co-ordinate = 10 sums up to 20 so that can be included in to the answer too.

OpenStudy (rushwr):

But in the given answers Z is paired up with j which is not greater than 4 so there one answers is wrong therefore we can't take that answer as the correct one :)

OpenStudy (rushwr):

Oops I didn't see the new question cumn. Sorry this is the explanation for the 2nd question u posted here

OpenStudy (rockycatsrule):

its ok

OpenStudy (rushwr):

can u open up a new question for that. I'll check that :)

OpenStudy (rockycatsrule):

yeah